OpenStudy (prepfortears):

Verify each trigonometric equation by substituting identities to math the right hand side of the equation to the left hand side of the equation (sin(x))/(1-cos(x))+(sin(x))/(1+cos(x))=2csx(x)

OpenStudy (prepfortears):

\[\frac{ \sin(x) }{ 1-\cos(x)}+\frac{ \sin(x) }{ 1+\cos(x)}=2 \csc(x)\]

OpenStudy (anonymous):

combine the terms on the left side using addition of fractions

OpenStudy (butterflydreamer):